OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. u = <-5, 8>, v = <-4, 8>

OpenStudy (solomonzelman):

\(\Large\color{black}{ \displaystyle \theta=\rm{ArcCosine}\left(\frac{\vec{\bf u}\cdot \vec{\bf v}}{\left|\left|\vec{\bf u}\right|\right|\cdot\left|\left|\vec{\bf v}\right|\right|}\right) }\)

OpenStudy (solomonzelman):

where, \(\left|\left|\vec{\bf v}\right|\right|\) demotes "the maginute of vector v"

OpenStudy (solomonzelman):

And, I said the exact same thing in your previous question. In any case, I hope that helps.

OpenStudy (anonymous):

im very confused